Please help me solve the following problem. REally appreciate the help.
Social Security benefits. If you earned an average of $25,000 over your working life and you retire after 2005 at age 62, 63, 64, then your annual Social Security benefit will be $7000,$7500, $8000, respectively (Social Security Administration. There is a liner equation that gives the annual benefit b in terms of age a for these three years. Find the equation

Answer by stanbon(75887)

Social Security benefits. If you earned an average of $25,000 over your working life and you retire after 2005 at age 62, 63, 64, then your annual Social Security benefit will be $7000,$7500, $8000, respectively (Social Security Administration. There is a liner equation that gives the annual benefit b in terms of age a for these three years. Find the equation
You have 3 points (62,7000),(63,7500), (64,8000)
Use any two to solve for m and b in the form y=mx+b
m=[8000-7000)/[64-62]=500
7000=500(62)+b
b=-24000
EQUATION:
Benefit=500(age)-24000
Cheers,
Stan H.